OpenStudy (anonymous):

Please help me solve this precalc problem: Find a function with given zeros 0, -3, and 4. Write the function in factored form and as a polynomial in standard form. Factored form:______ Standard form:________

OpenStudy (anonymous):

@Hero will you help me please?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (jdoe0001):

\(\bf \{x:0,-3,4\}\implies \begin{cases} x=0\to x-0=0\to &(x)=0\\ x=-3\to x+3=0\to &(x+3)=0\\ x=4\to x-4=0\to &(x-4)=0 \end{cases} \\ \quad \\ (x)(x+3)(x-4)=\textit{original polynomial}\impliedby \textit{factored form}\) notice, 3 factors, thus is called the "factored form" to make it in standard form multiply them all and arrange their term degree in descending order, from left to right http://www.mathsisfun.com/definitions/standard-form.html

OpenStudy (jdoe0001):

and make the equation equals to 0 or \(\bf (x)(x+3)(x-4)=0\)

OpenStudy (anonymous):

Ok I get it thank you!!! @jdoe0001

OpenStudy (jdoe0001):

yw